# Inference for Regression with

### v2

@@ 3.2.2 Sequence of Populations
An important insight developed in Theorem 2 is that $\kappa=0$ is in fact necessary for the validity of two-step inference. If $\kappa>0$, then the asymptotic distribution of $\hat{\bm{\psi}}$ has the correct variance (which is consistently estimated by the Eicker–Huber–White estimator computed using $\hat{\bm{\theta}}_{i}$) but its center is shifted away from the origin to $\kappa\bm{b}$ due to measurement error bias. [^12] This is the opposite of a generated regressors problem (Pagan, 1984), where the asymptotic variance is inflated but there is no location shift. With generated regressors there is a common finite-dimensional parameter estimated in the first stage whereas here all $n$ covariates $\bm{\theta}_{1},\ldots,\bm{\theta}_{n}$ are estimated in the first stage. See Bai and Ng, (2006) for further discussion in the context of factor-augmented regressions. Consequently, confidence intervals have the correct width but incorrect centering, and therefore have coverage below their nominal coverage. The bias, and hence the degree to which confidence intervals under-cover, is increasing in $\kappa$. It is worth noting that measurement error in $\hat{\bm{\theta}}_{i}$ will bias inference not just for $\bm{\gamma}$ but also for $\bm{\alpha}$ whenever $\bm{S}\bm{\theta}_{i}$ and $\mathbf{q}_{i}$ are correlated.
@@ 3.3.2 Similarity Measures
As $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ are not observed, a pragmatic two-step strategy is to estimate $\gamma_{1}$ by regressing $Y_{i}$ on
$(\hat{\mathbf{p}}_{1,i}\cdot\hat{\mathbf{p}}_{2,i}),$
where
@@
$b_{1}=-\left(\frac{\mathbb{E}[\mathbf{p}_{2,i}^{T}(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}\mathbf{p}_{1,i}^{T})\mathbf{p}_{2,i}]}{\mathrm{Var}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i})}\right)\gamma_{1},\quad b_{2}=-\left(\frac{\mathbb{E}[\mathbf{p}_{1,i}^{T}(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}\mathbf{p}_{2,i}^{T})\mathbf{p}_{1,i}]}{\mathrm{Var}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i})}\right)\gamma_{1}.$
As with our earlier results, Theorem 3 shows that $\hat{\mathbf{p}}_{1,i}$ and $\hat{\mathbf{p}}_{2,i}$ can be treated as if they are the true $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ for inference on $\gamma_{1}$ provided the amount of unstructured data per observation is “large” in the sense that both $\kappa_{1}=0$ and $\kappa_{2}=0$. For instance, in the patent example we would require both the amount of unstructured data per patent to be large (so that $\kappa_{1}=0$) and the amount of unstructured data for the existing stock to be large (so that $\kappa_{2}=0$). Otherwise, the location of two-step confidence intervals if shifted towards the origin which leads to biased inference.
@@ 4 One-Step Strategy
In this section, we first discuss from a theoretical perspective the one-step strategy for inference that overcomes the bias from the two-step strategy. While the idea is generic, for concreteness we develop it within a specific model we call the Supervised Topic Model with Covariates which extends the model presented in the previous section. Second, we discuss the computational challenge of implementing the one-step strategy, which we solve with Hamiltonian Monte Carlo (HMC) deployed on modern hardware systems. This allows for highly scalable inference with minimal coding. More in-depth overviews of HMC are provided in Neal, (2012), Hoffman and Gelman, (2014), and Betancourt, (2018). We are not aware of the application of HMC to topic models in the literature.
@@ 4.1 Supervised Topic Model with Covariates
The model for illustrating the one-step strategy again features the downstream regression model (4) and the upstream topic model (5). But we further enrich the model to include a probabilistic relationship between the topic shares $\bm{\theta}_{i}$ and a vector of $J$ covariates $\mathbf{g}_{i}$. The covariates $\mathbf{g}_{i}$ may or may not be the same as $\mathbf{q}_{i}$. We allow these additional dependencies to enhance the applicability of the model, and to demonstrate how our computational approach can perform inference for even complex models with relative ease.
Example: Monetary Policy Speeches (Continued).
To return to the example of Section 3, the downstream regression model (4) could capture how policymakers’ attention predicts policy actions controlling for economic conditions. But policymakers’ attention can itself be a function of speaker characteristics such as demographic variables, or past experience of economic conditions (Malmendier et al., 2021). Such variables would enter $\mathbf{g}_{i}$ but arguably not directly affect policy decisions beyond their effect on attention; i.e., they would not enter $\mathbf{q}_{i}$.
To capture dependence between $\bm{\theta}_{i}$ and $\mathbf{g}_{i}$, we specify the distribution of $\bm{\theta}_{i}$ conditional on $\mathbf{g}_{i}$ as logistic normal, though other specifications could also be used. The full model, which we call the Supervised Topic Model with Covariates (STMC), is specified in Model 1.
@@
The additional parameters in Model 1 are a $K\times J$ matrix of coefficients $\bm{\Phi}$ and scale parameters $\sigma_{\theta}$ and $\sigma_{Y}$. The $k$th row of $\bm{\Phi}$, denoted $\bm{\phi}_{k}$, captures how variation in covariates maps to variation in the prevalence of the $k$th topic across observations. Hence, a number of research questions can be addressed by performing inference on $\bm{\Phi}$. While we have modeled the error terms in the downstream regression and upstream logistic normal as homoskedastic to simplify presentation, this can easily be relaxed. Similarly, the logistic normal and normal specifications can also be substituted for other distributions or quasi-distributions as appropriate. For instance, one could use the sandwich quasi-likelihood of Müller, (2013).
To our knowledge, STMC is new in the literature. Roberts et al., (2014) presents a model in which a logistic normal distribution over $\bm{\theta}_{i}$ is parameterized by covariates but without a downstream regression. Blei and McAuliffe, (2010) and Ahrens et al., (2021) present models in which linear combinations of topic shares explain a response variable, but do not allow covariates to enter the distribution over $\bm{\theta}_{i}$. As such, we view STMC as of independent interest in the literature on topic modeling, although its primary purpose is to provide an example in which dimensionality reduction and linear regression are part of the same joint model and one cares about doing valid inference on model parameters.
4.2 Inference Approach for One-Step Strategy
The components of Model 1 combine to give a likelihood $l(\mathbf{x}_{i},Y_{i},\bm{\theta}_{i}|C_{i},\mathbf{g}_{i},\mathbf{q}_{i})$ for $\mathbf{x}_{i}$, $Y_{i}$, and $\bm{\theta}_{i}$ conditional on $C_{i}$ and covariates $\mathbf{g}_{i}$ and $\mathbf{q}_{i}$. As $\bm{\theta}_{i}$ is latent, we can integrate it out to obtain a likelihood $l(\mathbf{x}_{i},Y_{i}|C_{i},\mathbf{g}_{i},\mathbf{q}_{i})$ depending only on observable variables, which can then be used for maximum likelihood estimation of model parameters $\bm{\delta}=(\mathbf{B},\bm{\Phi},\bm{\gamma},\bm{\alpha},\sigma_{Y},\sigma_{\theta})$. However, there are two challenges. First, the integration has no closed-form solution and so must be performed numerically. Moreover, this numerical integration is high-dimensional and must be done observation-by-observation. As such, standard likelihood-based estimation is not computationally feasible.
The inference approach we take, while frequentist, is instead based on Bayesian computation. The integration step is performed implicitly as part of the sampling procedure. Similar approaches are taken to deal with latent states in Bayesian estimation of DSGE models (Herbst and Schorfheide, 2016). In this approach, Model 1 is supplemented with a prior for $\bm{\delta}$. The latent $\bm{\theta}_{i}$ are themselves treated as “parameters”, with the logistic normal component of Model 1 acting as their prior. We sample from the posterior distribution of $\left(\bm{\delta},(\bm{\theta}_{i})_{i=1}^{n}\right)$ conditional on the observed data $(\mathbf{x}_{i},Y_{i},C_{i},\mathbf{g}_{i},\mathbf{q}_{i})_{i=1}^{n}$. The marginal draws for $\bm{\delta}$ represent draws from the posterior distribution for $\bm{\delta}$ based on the integrated likelihood $l(\mathbf{x}_{i},Y_{i}|C_{i},\mathbf{g}_{i},\mathbf{q}_{i})$.
@@ 4.2 Inference Approach for One-Step Strategy
If one so desires, these priors can be changed with one line of code in our implementation of the inference algorithm explained below. In total, the model has eight hyperparameters: the three $\sigma$ terms in (Priors) as well as $\sigma^{2}_{\theta}$ in (Upstream Topic Model); the symmetric Dirichlet parameter $\eta$ in (Priors); the two Gamma distribution parameters in (Priors).
We emphasize again that the model and priors serve primarily as an illustration. An interested researcher should be able to modify it as needed to accommodate different data; to test robustness of the conclusions to specifying alternative distributions for the data; or to test robustness with respect to choice of priors. The key is to avoid having to re-derive complex inference algorithms every time the model is adjusted, and this is precisely the main advantage of automatic inference methods we now describe.
4.3 Overview of the HMC Algorithm
Our problem is to sample from the posterior distribution $q(\bm{\zeta}|(\mathbf{x}_{i},Y_{i},C_{i},\mathbf{g}_{i},\mathbf{q}_{i})_{i=1}^{n})$ where $\bm{\zeta}=\left(\bm{\delta},(\bm{\theta}_{i})_{i=1}^{n}\right)$ are the STMC parameters. To do so, we use Hamiltonian Monte Carlo (HMC), a modern Markov chain Monte Carlo (MCMC) algorithm that is particularly well-suited to high-dimensional models. [^13] Gibbs sampling is often used in the topic modeling literature (Griffiths and Steyvers, 2004). This is difficult to implement for STMC because (i) the logistic normal prior is not conjugate to the multinomial distribution and (ii) there are a large number of parameters in the model in realistic use cases. MCMC algorithms define a stochastic process, i.e., a Markov chain, whose ergodic distribution coincides with the posterior distribution one wishes to sample from. Samples from this Markov chain can be used to form estimates of interest, e.g. the expected value of a model parameter under the posterior distribution, as in Monte Carlo simulation. Efficient MCMC algorithms have low autocorrelation across samples which improves the accuracy of the resulting estimates.
4.4 HMC and Probabilistic Programming
Appendix C displays the NumPyro code needed to draw samples from the posterior distribution of STMC. The core code is only several dozen lines long, and individual elements can be quickly modified to specify alternative distributional assumptions or new models.
@@ 5 Empirical Results
where the error terms are drawn from normal distributions whose variances are assumed the same as in the one-step strategy. The prior distributions over the regression coefficients are also the same in both strategies. This procedure is designed to emulate the typical approach in the empirical literature while ensuring that any observed differences between the two strategies are not driven by different inference methods or implicit modeling choices.
@@ 5.1 Simulation
We start with a simulation exercise that compares the one- and two-step strategies in terms of (i) the evolution of the bias in regression coefficients across different values of $\kappa$ and (ii) the coverage of confidence intervals. We simulate the data according to the data generating process described in Model (1). [^16] We impose the anchor word assumptions in the simulation in the following way. We first draw $\tilde{\bm{\beta}_{1}}$ and $\tilde{\bm{\beta}_{2}}$ from symmetric $V-50$–dimensional Dirichlet priors. Then we insert 50 zeros into both $\tilde{\bm{\beta}_{1}}$ and $\tilde{\bm{\beta}_{2}}$ in such that there is no feature $v$ where $\beta_{v,1}=\beta_{v,2}=0$. Data is then simulated from these modified topic-feature distributions. We conduct three sets of simulations. Within each set, the amount of unstructured data per observation is the same for all observations and equal to $C_{i}=C\in\{10,25,200\}$. Together with the total number of observations, $n=10000$, this implies $\kappa\in\{10,4,0.5\}$, for the three sets of simulations, respectively. We conduct 200 simulations per set. Further details are included in Appendix B.
@@
Figure 1 presents the results. Each panel shows the coverage rates of confidence intervals for different parameter values: the share of simulations in which the values of the parameters are included in the 95% confidence intervals. The grey vertical dashed lines show the true value of the parameter. The blue vertical dashed line represents the median (across simulations) of mean posterior estimates for the two-step strategy. The two top panels show the results for the set of simulations where the amount of unstructured data per observation is the smallest and so $\kappa=10$ is relatively large. The theory in Section 3 suggests that in this case we should expect the two-step strategy to perform badly. This is indeed the case. The median (across simulations) estimate of $\gamma_{1}$ in top left, and $\phi_{1}$ in top right, are both substantially biased towards zero. Further, as predicted by theory, the width of the CIs using the two-step strategy is similar to the infeasible estimator that uses the true $\theta_{i}$. This, together with the bias, means that the CIs based on the two-step strategy under-cover. As reported in Table 1, for $\gamma_{1}$ the true value is included in the 95% CI in only 115/200 (57.5%) of simulations. For $\phi_{1}$ this looks even worse: the true value is never included in the CIs.
@@
On the other hand, the one-step strategy performs very well. The estimates appear unbiased, and the CIs have close to expected coverage. The coverage is $95.5$ and $92\%$ respectively for $\gamma_{1}$ and $\phi_{1}$. The difference between the lengths of CIs using the one-step strategy and those using the infeasible estimator is small but noticeable. In the former, $\theta_{i}$ is recognized as latent and the uncertainty in $\theta_{i}$ is accounted for when performing inference on $\gamma_{1}$ and $\phi_{1}$. The resulting CIs are approximately $20\%$ wider than those obtained with the infeasible estimator that uses the true $\theta_{i}$.

### v3

@@ 3.2.2 Sequence of Populations
An important insight developed in Theorem 2 is that $\kappa=0$ is in fact necessary for the validity of two-step inference. If $\kappa>0$, then the asymptotic distribution of $\hat{\bm{\psi}}$ has the correct variance (which is consistently estimated by the Eicker–Huber–White estimator computed using $\hat{\bm{\theta}}_{i}$) but its center is shifted away from the origin to $\kappa\bm{b}$ due to measurement error bias. [^10] This is the opposite of a generated regressors problem (Pagan, 1984), where the asymptotic variance is inflated but there is no location shift. With generated regressors there is a common finite-dimensional parameter estimated in the first stage whereas here all $n$ covariates $\bm{\theta}_{1},\ldots,\bm{\theta}_{n}$ are estimated in the first stage. See Bai and Ng, (2006) for further discussion in the context of factor-augmented regressions. Consequently, confidence intervals have the correct width but incorrect centering, and therefore have coverage below their nominal coverage. The bias, and hence the degree to which confidence intervals under-cover, is increasing in $\kappa$. It is worth noting that measurement error in $\hat{\bm{\theta}}_{i}$ will bias inference not just for $\bm{\gamma}$ but also for $\bm{\alpha}$ whenever $\bm{S}\bm{\theta}_{i}$ and $\mathbf{q}_{i}$ are correlated. The implication is that, even when unstructured data is used to create controls variables (see Avivi, (2024) for a recent example), the two-step strategy can bias inference.
@@ 3.3.2 Similarity Measures
As $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ are not observed, a pragmatic two-step strategy is to estimate $\gamma_{1}$ by regressing $Y_{i}$ on $(\hat{\mathbf{p}}_{1,i}\cdot\hat{\mathbf{p}}_{2,i})$ where
@@
$b_{1}=-\left(\frac{\mathbb{E}[\mathbf{p}_{2,i}^{T}(\mathrm{diag}(\mathbf{p}_{1,i})-\mathbf{p}_{1,i}^{\phantom{T}}\mathbf{p}_{1,i}^{T})\mathbf{p}_{2,i}^{\phantom{T}}]}{\mathrm{Var}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i})}\right)\gamma_{1},\quad b_{2}=-\left(\frac{\mathbb{E}[\mathbf{p}_{1,i}^{T}(\mathrm{diag}(\mathbf{p}_{2,i})-\mathbf{p}_{2,i}^{\phantom{T}}\mathbf{p}_{2,i}^{T})\mathbf{p}_{1,i}^{\phantom{T}}]}{\mathrm{Var}(\mathbf{p}_{1,i}\cdot\mathbf{p}_{2,i})}\right)\gamma_{1}.$
As with our earlier results, Theorem 3 shows that $\hat{\mathbf{p}}_{1,i}$ and $\hat{\mathbf{p}}_{2,i}$ can be treated as if they are the true $\mathbf{p}_{1,i}$ and $\mathbf{p}_{2,i}$ for inference on $\gamma_{1}$ provided the amount of unstructured data per observation is “large” in the sense that both $\kappa_{1}=0$ and $\kappa_{2}=0$. For instance, in the patent example we would require both the amount of unstructured data per patent to be large (so that $\kappa_{1}=0$) and the amount of unstructured data for the existing stock to be large (so that $\kappa_{2}=0$). Otherwise, the location of two-step confidence intervals is shifted towards the origin, which leads to biased inference.
@@ 4 One-Step Strategy
Having shown the bias inherent to the popular two-step approach for unstructured data analysis, we now develop an alternative one-step approach that allows for valid inference. The starting point is to note that the topic model (5) provides a likelihood for $\mathbf{x}_{i}$ conditional on $C_{i}$ and $\bm{\theta}_{i}$. We next build a joint likelihood for $(\mathbf{x}_{i},Y_{i},\bm{\theta}_{i})$ by specifying a parametric distribution for the regression errors in (4). Finally, in the spirit of correlated random effects estimation in panel data models, we also specify a distribution for the topic shares $\bm{\theta}_{i}$ conditional on covariates $\mathbf{g}_{i}$. The covariates $\mathbf{g}_{i}$ ($J$ in total) may or may not be the same as $\mathbf{q}_{i}$. In practice, one must specify particular distributions to complete the specification of the likelihood, but for now we keep the discussion general to highlight the broad applicability of the approach.
Together, these components combine to give a likelihood $l(\mathbf{x}_{i},Y_{i},\bm{\theta}_{i}|C_{i},\mathbf{g}_{i},\mathbf{q}_{i})$ for $\mathbf{x}_{i}$, $Y_{i}$, and $\bm{\theta}_{i}$ conditional on $C_{i}$ and covariates $\mathbf{g}_{i}$ and $\mathbf{q}_{i}$. As $\bm{\theta}_{i}$ is latent, we can integrate it out to obtain a likelihood $l(\mathbf{x}_{i},Y_{i}|C_{i},\mathbf{g}_{i},\mathbf{q}_{i})$ depending only on observable variables, which can then be used for maximum likelihood estimation of model parameters $\bm{\delta}$, consisting of $\mathbf{B}$, $\bm{\gamma}$, $\bm{\alpha}$, and any other parameters in the regression error and topic share distributions. However, there are two challenges. First, the integration has no closed-form solution and so must be performed numerically. Moreover, this numerical integration is high-dimensional and must be done observation-by-observation. As such, standard likelihood-based estimation is not computationally feasible. In the remainder of this section, we discuss how we overcome this challenge and present a specific model—the Supervised Topic Model with Covariates—that we use for the empirical results.
4.1 Inference Approach for One-Step Strategy
The inference approach we take, while frequentist, is instead based on Bayesian computation. The integration step is performed implicitly as part of the sampling procedure. Similar approaches are taken to deal with latent states in Bayesian estimation of DSGE models (Herbst and Schorfheide, 2016). In this approach, we introduce a prior for the model parameters $\bm{\delta}$ and treat the latent $\bm{\theta}_{i}$ as “parameters” drawn from a distribution that potentially depends on covariates $\mathbf{g}_{i}$ (as discussed above). We sample from the posterior distribution of $\left(\bm{\delta},(\bm{\theta}_{i})_{i=1}^{n}\right)$ conditional on the observed data $(\mathbf{x}_{i},Y_{i},C_{i},\mathbf{g}_{i},\mathbf{q}_{i})_{i=1}^{n}$. The marginal draws for $\bm{\delta}$ represent draws from the posterior distribution for $\bm{\delta}$ based on the integrated likelihood $l(\mathbf{x}_{i},Y_{i}|C_{i},\mathbf{g}_{i},\mathbf{q}_{i})$.
4.1.1 Hamiltonian Monte Carlo
Our problem is to sample from the posterior distribution $q(\bm{\zeta}|(\mathbf{x}_{i},Y_{i},C_{i},\mathbf{g}_{i},\mathbf{q}_{i})_{i=1}^{n})$ where $\bm{\zeta}=\left(\bm{\delta},(\bm{\theta}_{i})_{i=1}^{n}\right)$. To do so, we use Hamiltonian Monte Carlo (HMC), a modern Markov chain Monte Carlo (MCMC) algorithm that is particularly well-suited to high-dimensional models. [^11] More in-depth overviews of HMC are provided in Neal, (2012), Hoffman and Gelman, (2014), and Betancourt, (2018). We are not aware of the application of HMC to topic models in the literature. MCMC algorithms define a stochastic process, i.e., a Markov chain, whose ergodic distribution coincides with the posterior distribution one wishes to sample from. Samples from this Markov chain can be used to form estimates and credible sets/confidence intervals. Efficient MCMC algorithms have low autocorrelation across samples which improves the accuracy of the resulting estimates.
4.1.2 Implementation with probabilistic programming
@@ 4.2 Supervised Topic Model with Covariates
We now specify distributional assumptions to fully formulate a likelihood for empirical analysis. We assume the regression errors in (4) are normally distributed and that the distribution of $\bm{\theta}_{i}$ conditional on $\mathbf{g}_{i}$ is logistic normal. These assumptions are made for illustrative purposes and applied researchers may modify them as desired. The resulting model, which we call the Supervised Topic Model with Covariates (STMC), is formalized in Model 1.
@@
The matrix $\bm{\Phi}$ is a $K\times J$ matrix of regression coefficients. The $k$th row of $\bm{\Phi}$, denoted $\bm{\phi}_{k}$, captures how variation in covariates maps to variation in the prevalence of the $k$th topic across observations. Hence, a number of research questions can be addressed by performing inference on $\bm{\Phi}$. Model 1 also introduces scale parameters $\sigma_{\theta}$ and $\sigma_{Y}$. While we have modeled the error terms in the downstream regression and upstream logistic normal as homoskedastic to simplify presentation, this can easily be relaxed.
Example: Monetary Policy Speeches (Continued).
To return to the example of Section 3, the downstream regression model could capture how policymakers’ attention predicts policy actions controlling for economic conditions. But policymakers’ attention can itself be a function of speaker characteristics such as demographic variables, or past experience of economic conditions (Malmendier et al., 2021). Such variables would enter $\mathbf{g}_{i}$ but arguably not directly affect policy decisions beyond their effect on attention; i.e., they would not enter $\mathbf{q}_{i}$.
To our knowledge, STMC is new in the literature. Roberts et al., (2014) presents a model in which a logistic normal distribution over $\bm{\theta}_{i}$ is parameterized by covariates but without a downstream regression. Blei and McAuliffe, (2010) and Ahrens et al., (2021) present models in which linear combinations of topic shares explain a normally distributed response variable, but do not allow covariates to enter the distribution over $\bm{\theta}_{i}$. As such, we view STMC as of independent interest in the literature on topic modeling, although its primary purpose is to provide an example in which dimensionality reduction and linear regression are part of the same joint model and one cares about doing valid inference on model parameters.
@@
In total, the model has seven hyperparameters: the three $\sigma^{2}$ terms in (Priors) as well as $\sigma^{2}_{\theta}$ in (Upstream Topic Model); the symmetric Dirichlet parameter $\eta$ in (Priors); the two Gamma distribution parameters in (Priors).
Appendix C displays the NumPyro code needed to draw samples from the posterior distribution of STMC. The core code is only several dozen lines long, and individual elements can be quickly modified to specify alternative distributions or models. An interested researcher should be able to modify it as needed to accommodate different data; to test robustness of the conclusions to specifying alternative distributions for the data; or to test robustness with respect to choice of priors. The key is to avoid having to re-derive complex inference algorithms every time the model is adjusted, and this is precisely the main advantage of automatic inference methods.
@@ 5 Empirical Results
where the error terms are normal. The prior distributions over the regression coefficients are the same in both strategies. This procedure is designed to emulate the typical approach in the empirical literature while ensuring that any observed differences between the two strategies are not driven by different inference methods or implicit modeling choices.
@@ 5.1 Simulation
We start with a simulation exercise that compares the one- and two-step strategies in terms of (i) the evolution of the bias in regression coefficients across different values of $\kappa$ and (ii) the coverage of confidence intervals. We simulate the data according to the data generating process described in Model (1). [^14] We impose the anchor word assumptions in the simulation in the following way. We first draw ${\bm{\beta}_{1}}$ and ${\bm{\beta}_{2}}$ from symmetric $(V-50)$–dimensional Dirichlet priors. Then we insert 50 zeros into both ${\bm{\beta}_{1}}$ and ${\bm{\beta}_{2}}$ in such that there is no feature $v$ where $\beta_{v,1}=\beta_{v,2}=0$. Data is then simulated from these modified topic-feature distributions. We conduct three sets of simulations. Within each set, the amount of unstructured data per observation is the same for all observations and equal to $C_{i}=C\in\{10,25,200\}$. Together with the total number of observations, $n=10000$, this implies $\kappa\in\{10,4,0.5\}$, for the three sets of simulations, respectively. We conduct 200 simulations per set. Further details are included in Appendix B.
@@
Figure 1 presents the results. Each panel shows the coverage rates of confidence intervals for different parameter values: the share of simulations in which the values of the parameters are included in the 95% confidence intervals. The grey vertical dashed lines show the true value of the parameter. The blue vertical dashed line represents the median (across simulations) of mean posterior estimates for the two-step strategy. The two top panels show the results for the set of simulations where the amount of unstructured data per observation is the smallest and so $\kappa=10$ is relatively large. The theory presented earlier suggests that in this case we should expect the two-step strategy to perform badly. This is indeed the case. The median (across simulations) estimate of $\gamma_{1}$ in top left, and $\phi_{1}$ in top right, are both substantially biased towards zero. Further, as predicted by theory, the width of the CIs using the two-step strategy is similar to the infeasible estimator that uses the true $\theta_{i}$. This, together with the bias, means that the CIs based on the two-step strategy under-cover. As reported in Table 1, for $\gamma_{1}$ the true value is included in the 95% CI in only 115/200 (57.5%) of simulations. For $\phi_{1}$ this looks even worse: the true value is never included in the CIs.
@@
On the other hand, the one-step strategy performs very well. The estimates appear unbiased, and the CIs have close to nominal coverage. The coverage is $95.5$ and $92\%$ respectively for $\gamma_{1}$ and $\phi_{1}$. The difference between the lengths of CIs using the one-step strategy and those using the infeasible estimator is small but noticeable. In the former, $\theta_{i}$ is recognized as latent and the uncertainty in $\theta_{i}$ is accounted for when performing inference on $\gamma_{1}$ and $\phi_{1}$. The resulting CIs are approximately $20\%$ wider than those obtained with the infeasible estimator that uses the true $\theta_{i}$.